AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 565     Entry time: Fri May 31 06:01:55 2024
Author: JM, TY, PW, EM, CB 
Subject: Run 12 - beam on target 

07:00

Run 12 started. D+ beam is incident on the target at 1.5 MeV

Beam intensity in the ring is ~4e7 particles

Target density shows some fluctuations but is ~1e11 atoms/cm2. Target temp 35K

Detectors are positioned fully out (Left arm 18.86mm and right arm 19.8 mm ) and are not moving

Ring pressure looks OK, YR09 E target sensor seems unaffected by the target currently reads ~5e-11 mbar.

Beam is in the ring for ~15s.

Temps and system checks OK

Leakage current OK

Merger and tape server OK

Horizontal beam position 8 mm and vertical beam position 0mm for overlap with target.

Statistics - attachment 1

Energy histogram - attachment 2

XY plot - attachment 3

We see high rate at 1.5 MeV in left dssds but very little in right dssds. XY plots also show counts are concerntrated in a very circular distribution.

07:15

Gleb has just called and requires access to the cave to check on a turbo pump in the source.

Ring vacuum looks good

Beam will be down for 10-15 minutes only hopefully. 

07:45

Beam is back in the ring but injection requires optimisation.

09:00

There are some issues with the source that require a few hours to resolve. Potentially until 7 or 8 pm tonight. We have about an hour before work starts.

We have done a quick test of the motors and moved the left and right arms in using the pnuematic only.

New position left arm = 62.71 mm , right arm = 64.6 mm.

We observed the rates in left detectors went up significantly. In the right detectors we observed little increase.

See attached xy plot - attachment 4

See attached energy histogram - attachment 5

Beam intensity has dropped significantly - below 1e6 particles so even at in positons we see few counts at the moment.

09:25

Source has degraded pretty rapidly so beam will be stopped for now.

Detectors moved back out.

Target nozzle moved from X= -13000, Y=2500 to X= -11000, Y=2500. S1 has dropped to 5e-10 mbar.

Note Gleb mentioned that last night it was likely a large proportion of the ions stored were molecular hydrogen not deuterium.

16:18

Temps and system checks OK

Leakage current OK

Merger and tape server OK

 

18:10

Temps and system checks OK

Leakage current OK

Merger and tape server OK

 

19:38

Operators are trying to optimize some beam. For now the beam is lower than it should be probably because is easy to work with the source like that.

Temps and system checks OK

Leakage current OK

Merger and tape server OK


 

Attachment 1: Screenshot_from_2024-05-31_07-32-33.png  117 kB  Uploaded Fri May 31 07:35:54 2024  | Hide | Hide all
Screenshot_from_2024-05-31_07-32-33.png
Attachment 2: Screenshot_from_2024-05-31_07-33-19.png  290 kB  Uploaded Fri May 31 07:35:54 2024  | Hide | Hide all
Screenshot_from_2024-05-31_07-33-19.png
Attachment 3: Screenshot_from_2024-05-31_07-33-57.png  532 kB  Uploaded Fri May 31 07:35:54 2024  | Hide | Hide all
Screenshot_from_2024-05-31_07-33-57.png
Attachment 4: Screenshot_from_2024-05-31_09-17-08.png  486 kB  Uploaded Fri May 31 09:18:51 2024  | Hide | Hide all
Screenshot_from_2024-05-31_09-17-08.png
Attachment 5: Screenshot_from_2024-05-31_09-21-40.png  287 kB  Uploaded Fri May 31 09:21:51 2024  | Hide | Hide all
Screenshot_from_2024-05-31_09-21-40.png
Attachment 6: Screenshot_from_2024-05-31_16-19-29.png  1.001 MB  Uploaded Fri May 31 16:28:02 2024  | Hide | Hide all
Screenshot_from_2024-05-31_16-19-29.png
Attachment 7: Screenshot_from_2024-05-31_16-29-22.png  975 kB  Uploaded Fri May 31 16:29:32 2024  | Hide | Hide all
Screenshot_from_2024-05-31_16-29-22.png
Attachment 8: Screenshot_from_2024-05-31_16-31-36.png  1.054 MB  Uploaded Fri May 31 16:31:46 2024  | Hide | Hide all
Screenshot_from_2024-05-31_16-31-36.png
Attachment 9: Screenshot_from_2024-05-31_18-12-36.png  1.011 MB  Uploaded Fri May 31 18:12:47 2024  | Hide | Hide all
Screenshot_from_2024-05-31_18-12-36.png
Attachment 10: Screenshot_from_2024-05-31_19-42-42.png  996 kB  Uploaded Fri May 31 19:45:32 2024  | Hide | Hide all
Screenshot_from_2024-05-31_19-42-42.png
Attachment 11: Screenshot_from_2024-05-31_19-45-55.png  581 kB  Uploaded Fri May 31 19:46:05 2024  | Hide | Hide all
Screenshot_from_2024-05-31_19-45-55.png
Attachment 12: Screenshot_from_2024-05-31_20-04-54.png  593 kB  Uploaded Fri May 31 20:05:30 2024  | Hide | Hide all
Screenshot_from_2024-05-31_20-04-54.png
ELOG V3.1.4-unknown